OpenStudy (stephanieelizzz):

Simplify: 2log3(3x) - log3(x2) --> I got it to log3(3x^2) - log3(x^2) but now what?

OpenStudy (sunnnystrong):

Okay so: Basic log properties --> 1.) \[\log_{a}x-\log_{a}y=\log_{a}\frac{ x }{ y } \] \[2\log_{3}(3x)-\log_{3}(x^2) \] = \[2\log_{3}\frac{ 3x }{ x^2 } \]

OpenStudy (sunnnystrong):

@stephanieelizzz ... how could you simplify this further?

jimthompson5910 (jim_thompson5910):

careful @stephanieelizzz you need to square ALL of 3x. Not just the x it should be (3x)^2 = 9x^2 instead of just 3x^2

jimthompson5910 (jim_thompson5910):

@sunnnystrong you need to move the 2 up into the exponent before you can use that rule

OpenStudy (sunnnystrong):

@jim_thompson5910 ooops okay haha \[\log_{3}\frac{ 9x^2 }{ x^2 }\]

OpenStudy (stephanieelizzz):

@jim_thompson5910 oopsiesssss thank you!

OpenStudy (stephanieelizzz):

Can you cancel out the x^2? @sunnnystrong

OpenStudy (sunnnystrong):

@jim_thompson5910 thank you! & Yes I would think soo hmmm not 100%. What do you think @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

yes the x^2 terms will cancel. Then you need to evaluate \(\Large \log_3(9)\)

OpenStudy (stephanieelizzz):

You'd end up with 2 after everything is said and done, I believe. :)

jimthompson5910 (jim_thompson5910):

yep since 9 = 3^2